Green Corporation, a closely held operating corporation, reports the following:

Taxable income                                                       $100,000

Long-term capital gain                                              130,000

Dividends-received deduction                                   120,000

Federal income taxes on long-term capital gain       compute

Accumulated earnings credit                                     120,000

Federal income taxes                                                compute

Calculate Green's accumulated taxable income
Answer #1

SOLUTION:

Taxable income 100,000
Plus: dividends - received deduction 120,000
Minus: LTCG (net of taxes)* -73,200
120,000 * 0.39% = 46,800; 120,000 - 46,800 = 73,200
Minus: Accumulated earnings credit -120,000
Minus: Federal income taxes -22,150
Accumulated taxable income 4,650
*120,000 * 0.39% = 46,800; 120,000 - 46,800 = 73,200
**Tax on first $100,000 of taxable income: 22,150